# via Mixture-of-Curvature Experts

### v1

@@ B.1 MiCE as a Lorentzian Module
In this section, we expand on the fact that MiCE is indeed a valid hyperbolic module throughout. Note that since Equation 6 consists of the combination of a Lorentzian residual connection [23] and Lorentzian centroid [29], it suffices to show that the projection from input manifold to expert manifold, and the reverse projection, are valid projections between Lorentz hyperbolic spaces. In fact, it suffices to show that given $\mathbf{x}\in\mathbb{L}^{K_{1},n}$, we have $\sqrt{K_{1}/K_{2}}\mathbf{x}\in\mathbb{L}^{K_{2},n}$. To see this, note that
@@
Thus, $\sqrt{K_{1}/K_{2}}\mathbf{x}\in\mathbb{L}^{K_{2},n}$ as desired. Then, each projection via scaling by $\sqrt{K/K_{s,i}}$ and $\sqrt{K/K_{r,i}}$ indeed map the input vector $\mathbf{x}$ to the expert manifold, and the projection via $\sqrt{K_{s,i}/K}$ and $\sqrt{K_{r,i}/K}$ maps the output of the experts back to the input manifold. As a result, every vector in Equation 6 lives on the input manifold, hence the output lives in $\mathbb{L}^{K,n}$ as desired.
Additionally, note that since the squared Lorentzian distance is given by $d^{2}_{\mathcal{L}}(\mathbf{x},\mathbf{y})=2/K-2\langle\mathbf{x},\mathbf{y}\rangle=2/K+2x_{t}y_{t}-2\mathbf{x}_{s}^{\top}\mathbf{y}_{s}$, it scales inversely w.r.t. $\mathbf{x}_{s}^{\top}\mathbf{y}_{s}$. As a result, the gating score obtained through Equation 5 is minimizing the the squared hyperbolic distance between the input token vector $\mathbf{x}_{t}$ and the vector $\mathbf{y}_{j}$ (by viewing centroid vector $\mathbf{y}_{j}$ as the space-like dimension of a Lorentz hyperbolic vector). Therefore, the gating module is in fact a hyperbolic module as well.
@@ B.2 Hyperbolic Multi-Head Latent Attention
Following previous works [11, 10], as RoPE is incompatible with MLA due to position coupling, we employ a decoupled HoPE scheme with HMLA, where we use additional query vectors with a shared key. Let $\mathbf{W^{QR}}\in\mathbb{R}^{(n_{q}+1)\times(h_{n}n_{r})}$ and $\mathbf{W^{KR}}\in\mathbb{R}^{(n_{kv}+1)\times n_{r}}$ be the upward projection matrix of the decoupled queries and the shared key respectively, where $n_{r}$ is the dimension per head. We apply HoPE to these vectors to obtain the position-encoded vectors
@@
where $\mathrm{HCat}$ denotes hyperbolic concatenation [48, 37]. The attention score is computed based on negative squared Lorentz distance similar to Equation 3 as
@@
where $\mathbf{o}_{t}=[\mathbf{o}_{t,1},\ldots,\mathbf{o}_{t,h_{n}}]$ and $\mathbf{W^{O}}\in\mathbb{R}^{h_{n}(n+1)\times h_{n}n}$ is the out-project matrix. HMLA enables HELM models to improve computational efficiency during training and inference compared to the regular hyperbolic self-attention in Equation 3.
@@ B.3 Hyperbolic SwiGLU Feedforward Network
In this section, we introduce hyperbolic SwiGLU feedforward networks (FFNs), whose Euclidean formulation is widely used in LLMs [18, 10]. This differs from previous FNNs used in hyperbolic Transformers in the need for feature multiplication and activation function [5, 6, 48]. Let $\mathbf{x}\in\mathbb{L}^{K,n}$ be the input tokens, $\mathbf{W}_{1},\mathbf{W}_{3}\in\mathbb{R}^{(n+1)\times m}$ be the weights of internal projection layers and $\mathbf{W}_{2}\in\mathbb{R}^{(m+1)\times n}$ be the weight matrix of the outward projection layer. Then, the hyperbolic SwiGLU FNN $\mathrm{HFFN}_{SG}:\mathbb{L}^{K,n}\to\mathbb{L}^{K,n}$ is given by
@@
where $\mathrm{SiLU}_{\mathcal{L}}$ denotes SiLU activation using the HRC activation operations from Hypformer [48] and $\otimes_{s}$ denotes multiplication on the space dimention of a Lorentz vector, i.e. $\mathbf{x}\otimes_{s}\mathbf{y}=\mathbf{x}_{s}\mathbf{y}_{s}$.
@@ C.1 Models Setup
HELM-MiCE model setup. We follow the notation in Section B.2 and Section 4.2. For HELM-MiCE, we used HMLA as the attention mechanism in the attention block, MiCE as the sparse feedforward network, and $\mathrm{HFNN}_{SG}$ as the dense feedforward network. For both sizes, only the first decoder block uses the dense layer and the rest of the blocks using the MiCE layer. The MiCE layers use $\mathrm{HFNN}_{SG}$ as well for its feedforward component. For the dense layer, we set the intermediate dimension to be $4h_{n}n$. For the MiCE layers, we set the intermediate dimension of $\mathrm{HFNN}_{SG}$ as $2h_{n}n$.
For the $\sim 100M$ sized model, we used 6 total layers, 6 heads ($n_{h}=6$) each with $n=64$, and we set $n_{kv}=64$, $n_{r}=16$. For MiCE layers, we employ 4 experts with 2 active experts per token ($N_{r}=4,K_{r}=2$). We use one shared expert ($N_{s}=1$). For the curvatures of the routed experts, we set them to be uniform from $-0.1$ to $2.0$. The curvature of the shared expert is set to be $-1$. The curvature of the entire model is set to $-1$ as well. For HMLA layers, in practice, the upward projection matrices do not need to project back to the full dimension of $h_{n}n$. Due to compute constraints, we instead employ a reduction in dimensionality during the upward projection, where $\mathbf{W^{UK}},\mathbf{W^{UV}}\in\mathbb{R}^{(n_{kv}+1)\times h_{n}n/2}$, and the outward projection matrix $\mathbf{W^{O}}$ projects back to the full dimensionality of the input with $\mathbf{W^{O}}\in\mathbb{R}^{h_{n}(n/2+1)\times h_{n}n}$.
For the $\sim 1B$ sized model, we use 16 total layers, 14 heads ($n_{h}=14$) each with $n=64$, and we set $n_{kv}=256$, $n_{r}=64$. For MiCE layers, we employ 4 experts with 2 active experts per token ($N_{r}=8,K_{r}=2$). We use one shared expert ($N_{s}=1$). For the curvatures of the routed experts, we set them to be uniform from $-0.1$ to $2.0$. The curvature of the shared expert is set to be $-1$. The curvature of the entire model is set to $-1$ as well. We do not use the same reduction in dimensionality during upward projection as we did in the $\sim 100M$ case to enable for more expressive attention modules.
HELM-D model setup. For the HELM-D model, we only train the ${}100M$ sized model. Here, we use 6 layers, 6 heads each with dimension 64, and we set the intermediate dimension of the $\mathrm{HFNN}_{SG}$ feedforward networks to be 4 times the total model dimension. We set the overall curvature of the model to $-1$. All hyperbolic models are built on top of HyperCore He et al. [22].
@@ C.2 Training Details
Dataset. For the training dataset, we use the English portion of the Wikipedia dataset [14]. This dataset consists of $\sim 6.4M$ rows of data. We download the dataset directly from Huggingface. The raw text data is then passed through the LLaMA3.1-8B tokenizer [18], which has a vocabulary size of $\sim 128$K. We use a sequence length of $2048$ for all models. Samples longer than $2048$ tokens were broken up into multiple samples, with the trailing tailed dropped. The tokenized dataset consist of roughly $4.5B\sim 5B$ tokens. For training efficiency, as we measured the average number of tokens per sample is $\sim 700$ across the dataset, we used sample packing with a packing ratio of $3.0$. Then packed samples shorted than $2048$ tokens are then padded on the right.
Pipeline setup. For training, we set up data-parallelism with Hugginface Accelerate. We use an effective batch size of $\sim 2M$ tokens (including padding). To ensure a fair comparison between the hyperbolic and Euclidean models, we use a learning rate of 2e-4 for all dense models and a learning rate of 4e-4 for the MoE and MiCE models. A weight decay rate of 0.01 was used for all models. For the HELM-MiCE models and the DeepSeek models, in order to balance the load between each expert, we utilize the auxiliary-loss-free load balancing strategy and the complementary sequence-wise auxiliary loss during training. The former punishes extreme load imbalance among the experts by dynamically updating a bias term during the gating module, while not needing an explicit auxiliary loss computation for better training efficiency. The latter punishes extreme load imbalance for any particular sequence. All training used a cosine annealing learning rate scheduler with a final target learning rate of 0.1×0.1$\times$0.1 × the initial learning rate, with $3\%$ of the gradient update steps used as warmup steps.
@@ C.3 Evaluation Details
Appendix D Ablation Studies
In this section, we perform additional ablation studies to access the effectiveness of HoPE and HMLA.
D.1 Ablation for HMLA
@@
[…]
HELM-MiCE 120M $\mathbf{19.3}$ $\mathbf{26.2}$ $\mathbf{27.4}$ $\mathbf{24.7}$ $\underline{24.1}$ $\mathbf{24.2}$
@@
HELM-D 115M $\underline{20.1}$ $\underline{25.9}$ $27.0$ 25.8 $21.2$ 24.0
[…]
@@
HELM-MiCE 120M $19.3$ 26.0 $\underline{27.4}$ $\underline{24.7}$ $\underline{23.5}$ 24.2
Past works have found that in the Euclidean case, Multi-Head Latent Attention can achieve comparable and in some cases even superior performance compared to regular Multi-Head Attention [10]. Here we assess the effectiveness of HMLA against hyperbolic Multi-Head self-Attention. We train HELM-MiCE with the same setup, where we replace the HMLA layers with a hyperbolic Multi-Head self-Attention layer as given in Equation 3. We denote the this model as MiCE-HMHA. The results are shown in Table 3. HELM-MiCE outperforms MiCE-HMHA in 3 out of the 5 tasks, achieving the same accuracy for 1 task, with the MiCE-HMHA achieving better accuracy in the last task. The results demonstrate the effectiveness of HELM-MiCE while significantly reducing the memory footprint of the KV-cache. Both hyperbolic models still outperform the Euclidean model, demonstrating the effectiveness of HELM in general.
D.2 Ablation for HoPE
In this section we assess the effectiveness of HoPE against other hyperbolic positional encoding methods, namely the learned relative positional encoding from Hypformer [48]. We devise a variant of HELM-D, denoted as HELM-D-L and a variant of HELM-MiCE, denoted as HELM-MiCE-L, where each model uses the learned positional encoding instead of HoPE. The results are shown in Table 4. Overall both HELM-MiCE and HELM-D outperform their counterparts that use learned positional encoding instead of HoPE. Interestly, however, HELM-MiCE-L and HELM-D-L outperformed HELM-MiCE and HELM-D respectively on the ARC-Challenging benchmark, possibly due to better alignment with reasoning prompts with non-uniform encodings. Nevertheless, the results demonstrate the effectiveness of HoPE over learned positional encodings in 4 out of the 5 tasks.

### v2

@@ B.1 MiCE as a Lorentzian Module
In this section, we expand on the fact that MiCE is indeed a valid hyperbolic module throughout. Note that since Equation˜6 consists of the combination of a Lorentzian residual connection [he2025lresnet] and Lorentzian centroid [law2019lorentzian], it suffices to show that the projection from input manifold to expert manifold, and the reverse projection, are valid projections between Lorentz hyperbolic spaces. In fact, it suffices to show that given $\mathbf{x}\in\mathbb{L}^{K_{1},n}$, we have $\sqrt{K_{1}/K_{2}}\mathbf{x}\in\mathbb{L}^{K_{2},n}$. To see this, note that
@@
Thus, $\sqrt{K_{1}/K_{2}}\mathbf{x}\in\mathbb{L}^{K_{2},n}$ as desired. Then, each projection via scaling by $\sqrt{K/K_{s,i}}$ and $\sqrt{K/K_{r,i}}$ indeed map the input vector $\mathbf{x}$ to the expert manifold, and the projection via $\sqrt{K_{s,i}/K}$ and $\sqrt{K_{r,i}/K}$ maps the output of the experts back to the input manifold. As a result, every vector in Equation˜6 lives on the input manifold, hence the output lives in $\mathbb{L}^{K,n}$ as desired.
Additionally, note that since the squared Lorentzian distance is given by $d^{2}_{\mathcal{L}}(\mathbf{x},\mathbf{y})=2/K-2\langle\mathbf{x},\mathbf{y}\rangle=2/K+2x_{t}y_{t}-2\mathbf{x}_{s}^{\top}\mathbf{y}_{s}$, it scales inversely w.r.t. $\mathbf{x}_{s}^{\top}\mathbf{y}_{s}$. As a result, the gating score obtained through Equation˜5 is minimizing the the squared hyperbolic distance between the input token vector $\mathbf{x}_{t}$ and the vector $\mathbf{y}_{j}$ (by viewing centroid vector $\mathbf{y}_{j}$ as the space-like dimension of a Lorentz hyperbolic vector). Therefore, the gating module is in fact a hyperbolic module as well.
@@ B.2 Hyperbolic Multi-Head Latent Attention
Following previous works [deepseekv2, deepseekai2024deepseekv3technicalreport], as RoPE is incompatible with MLA due to position coupling, we employ a decoupled HoPE scheme with HMLA, where we use additional query vectors with a shared key. Let $\mathbf{W^{QR}}\in\mathbb{R}^{(n_{q}+1)\times(h_{n}n_{r})}$ and $\mathbf{W^{KR}}\in\mathbb{R}^{(n_{kv}+1)\times n_{r}}$ be the upward projection matrix of the decoupled queries and the shared key respectively, where $n_{r}$ is the dimension per head. We apply HoPE to these vectors to obtain the position-encoded vectors
@@
where $\mathrm{HCat}$ denotes hyperbolic concatenation [yang2024hypformer, qu2022autoencoding]. The attention score is computed based on negative squared Lorentz distance similar to Equation˜3 as
@@
where $\mathbf{o}_{t}=[\mathbf{o}_{t,1},\ldots,\mathbf{o}_{t,h_{n}}]$ and $\mathbf{W^{O}}\in\mathbb{R}^{h_{n}(n+1)\times h_{n}n}$ is the out-project matrix. HMLA enables HELM models to improve computational efficiency during training and inference compared to the regular hyperbolic self-attention in Equation˜3.
@@ B.3 Hyperbolic SwiGLU Feedforward Network
In this section, we introduce hyperbolic SwiGLU feedforward networks (FFNs), whose Euclidean formulation is widely used in LLMs [dubey2024llama3, deepseekai2024deepseekv3technicalreport]. This differs from previous FNNs used in hyperbolic Transformers in the need for feature multiplication and activation function [chen2021fully, chen2024hyperbolic, yang2024hypformer]. Let $\mathbf{x}\in\mathbb{L}^{K,n}$ be the input tokens, $\mathbf{W}_{1},\mathbf{W}_{3}\in\mathbb{R}^{(n+1)\times m}$ be the weights of internal projection layers and $\mathbf{W}_{2}\in\mathbb{R}^{(m+1)\times n}$ be the weight matrix of the outward projection layer. Then, the hyperbolic SwiGLU FNN $\mathrm{HFFN}_{SG}:\mathbb{L}^{K,n}\to\mathbb{L}^{K,n}$ is given by
@@
where $\mathrm{SiLU}_{\mathcal{L}}$ denotes SiLU activation using the HRC activation operations from Hypformer [yang2024hypformer] and $\otimes_{s}$ denotes multiplication on the space dimention of a Lorentz vector, i.e. $\mathbf{x}\otimes_{s}\mathbf{y}=\mathbf{x}_{s}\mathbf{y}_{s}$.
@@ C.1 Models Setup
HELM-MiCE model setup. We follow the notation in Section˜B.2 and Section˜4.2. For HELM-MiCE, we used HMLA as the attention mechanism in the attention block, MiCE as the sparse feedforward network, and $\mathrm{HFNN}_{SG}$ as the dense feedforward network. For both sizes, only the first decoder block uses the dense layer and the rest of the blocks using the MiCE layer. The MiCE layers use $\mathrm{HFNN}_{SG}$ as well for its feedforward component. For the dense layer, we set the intermediate dimension to be $4h_{n}n$. For the MiCE layers, we set the intermediate dimension of $\mathrm{HFNN}_{SG}$ as $2h_{n}n$.
For the $\sim 100M$ sized model, we used 6 total layers, 6 heads ($n_{h}=6$) each with $n=64$, and we set $n_{kv}=64$, $n_{r}=16$. For MiCE layers, we employ 4 experts with 2 active experts per token ($N_{r}=4,K_{r}=2$). We use one shared expert ($N_{s}=1$). For the curvatures of the routed experts, we set them to be uniform from 0.1-0.1 to $2.0$. The curvature of the shared expert is set to be 1-1. The curvature of the entire model is set to 1-1 as well. For HMLA layers, in practice, the upward projection matrices do not need to project back to the full dimension of $h_{n}n$. Due to compute constraints, we instead employ a reduction in dimensionality during the upward projection, where $\mathbf{W^{UK}},\mathbf{W^{UV}}\in\mathbb{R}^{(n_{kv}+1)\times h_{n}n/2}$, and the outward projection matrix $\mathbf{W^{O}}$ projects back to the full dimensionality of the input with $\mathbf{W^{O}}\in\mathbb{R}^{h_{n}(n/2+1)\times h_{n}n}$.
For the $\sim 1B$ sized model, we use 16 total layers, 14 heads ($n_{h}=14$) each with $n=64$, and we set $n_{kv}=256$, $n_{r}=64$. For MiCE layers, we employ 4 experts with 2 active experts per token ($N_{r}=8,K_{r}=2$). We use one shared expert ($N_{s}=1$). For the curvatures of the routed experts, we set them to be uniform from 0.1-0.1 to $2.0$. The curvature of the shared expert is set to be 1-1. The curvature of the entire model is set to 1-1 as well. We do not use the same reduction in dimensionality during upward projection as we did in the $\sim 100M$ case to enable for more expressive attention modules.
HELM-D model setup. For the HELM-D model, we only train the 100M~100M sized model. Here, we use 6 layers, 6 heads each with dimension 64, and we set the intermediate dimension of the $\mathrm{HFNN}_{SG}$ feedforward networks to be 4 times the total model dimension. We set the overall curvature of the model to 1-1. All hyperbolic models are built on top of HyperCore he2025hypercore.
@@ C.2 Training Details
Dataset. For the training dataset, we use the English portion of the Wikipedia dataset [wikidump]. This dataset consists of $\sim 6.4M$ rows of data. We download the dataset directly from Huggingface. The raw text data is then passed through the LLaMA3.1-8B tokenizer [dubey2024llama3], which has a vocabulary size of $\sim 128$K. We use a sequence length of $2048$ for all models. Samples longer than $2048$ tokens were broken up into multiple samples, with the trailing tailed dropped. The tokenized dataset consist of roughly $4.5B\sim 5B$ tokens. For training efficiency, as we measured the average number of tokens per sample is $\sim 700$ across the dataset, we used sample packing with a packing ratio of $3.0$. Then packed samples shorted than $2048$ tokens are then padded on the right.
Pipeline setup. For training, we set up data-parallelism with Hugginface Accelerate. We use an effective batch size of $\sim 2M$ tokens (including padding). To ensure a fair comparison between the hyperbolic and Euclidean models, we use a learning rate of 2e-4 for all dense models and a learning rate of 4e-4 for the MoE and MiCE models. A weight decay rate of 0.01 was used for all models. For the HELM-MiCE models and the DeepSeek models, in order to balance the load between each expert, we utilize the auxiliary-loss-free load balancing strategy and the complementary sequence-wise auxiliary loss during training. The former punishes extreme load imbalance among the experts by dynamically updating a bias term during the gating module, while not needing an explicit auxiliary loss computation for better training efficiency. The latter punishes extreme load imbalance for any particular sequence. All training used a cosine annealing learning rate scheduler with a final target learning rate of $0.1\times$ the initial learning rate, with $3\%$ of the gradient update steps used as warmup steps.
@@ C.3 Evaluation Details
Appendix D Ablation Studies and Additional Experiments
In this section, we perform additional experiments such as computational cost analysis, ablation studies to access the effectiveness of HoPE and HMLA, and comparisons with prior works of hyperbolic language models.
Table 4: Runtime and peak memory usage per iteration comparison between HELM variants and Euclidean counterparts. HELM models are within 1.55X in runtime and 1.11X in memory usage of the Euclidean counterparts for both the 100M and 1B models.
Model # Params Runtime Memory Usage
LLaMA 100M $8.4$s $23.8$ GB
HELM-D 100M $13.1$s $24.9$ GB
DeepSeekV3 100M $11.0$s $23.5$ GB
HELM-MiCE 100M $16.9$s $26.3$ GB
DeepSeekV3 1B $83.5$s $33.1$ GB
HELM-MiCE 1B $119.4$s $35.8$ GB
Table 5: Performance comparison between a 115M Hypformer model and HELM-D, where HELM-D consistently outperforms the baseline.
0-Shot 0-Shot 0-Shot 5-Shot 5-Shot -
Hypformer 115M $19.4$ $25.1$ $26.6$ $22.9$ $\mathbf{23.6}$ $23.5$
HELM-D 115M $\mathbf{20.3}\pm 0.2$ $\mathbf{25.9}\pm 0.1$ $\mathbf{27.1}\pm 0.4$ $\mathbf{25.6}\pm 0.2$ $21.4\pm 0.3$ $\mathbf{24.1}\pm 0.1$
D.1 Runtime and memory usage analysis
While HELM inevitably introduces computational overhead from operations that respect the curvature of the embedding space, our proposed methods are efficient in both runtime and memory usage. HELM models are within 1.55X in runtime and 1.11X in memory usage of the Euclidean counterparts for both the 100M and 1B models. In Table˜4, we show that runtime and peak memory usage comparisons between HELM and the Euclidean baselines for one training iteration (roughly 2M tokens). The results shown are for our exact experimental setup ran on 4 A100 GPUs, where we show the worst runtime and memory usage across the ranks. The results are averaged over 10 runs, and standard deviations are not shown since they are within 0.2 of the results.
D.2 Comparison with Prior Hyperbolic Language Models
Table 6: Machine translation BLEU score. HELM-D outperforms prior works of hyperbolic Transformers.
Model IWSLT’14 WMT’14
HAT gulcehre2019hyperbolicAT $23.7$ $21.8$
HNN++ HNN++ $22.0$ $25.5$
HyboNet chen2021fully $25.9$ $26.2$
HELM-D $\mathbf{26.3}$ $\mathbf{26.5}$
We performance additional experiments to compare the HELM architecture against prior works of hyperbolic language models and Transformers. As the majority of prior hyperbolic Transformers lacked essential components common in modern LLMs such as LayerNorm or positional encoding, we train a small version of HELM-D and compare the performance against these models in the machine translation task. The setup of our experiment is identical to that of chen2021fully. Hypformer yang2024hypformer, on the other hand, does possess all components. As a result, we compare the performance of HELM-D with Hypformer by training a 100M model on the same setup.
D.3 Ablation for HMLA
@@
MiCE-HMHA 120M $\underline{19.3}$ $\underline{25.7}$ $\underline{26.0}$ $23.8$ $\mathbf{25.3}$ $\underline{23.7}$
textbfHELM-MiCE 120M $\mathbf{19.7}\pm 0.3$ $\mathbf{25.9}\pm 0.2$} $\mathbf{27.7}\pm 0.4$ $\mathbf{24.4}\pm 0.2$ $\underline{23.2}\pm 0.5$ $\mathbf{24.1}\pm 0.1$
@@
HELM-D 115M $\underline{20.3\pm 0.2}$ $\mathbf{25.9}\pm 0.1$ $27.1\pm 0.4$ $\mathbf{25.6}\pm 0.2$ $21.4\pm 0.3$ $\mathbf{24.1}\pm 0.1$
DeepSeekV3 120M $19.2$ $25.2$ $23.4$ $24.2$ $21.8$ $22.8$
@@
HELM-MiCE 120M $19.7\pm 0.3$} $\mathbf{25.9}\pm 0.2$} $\underline{27.7}\pm 0.4$ $\underline{24.4}\pm 0.2$ $\underline{23.2}\pm 0.5$ $\mathbf{24.1}\pm 0.1$
Past works have found that in the Euclidean case, Multi-Head Latent Attention can achieve comparable and in some cases even superior performance compared to regular Multi-Head Attention [deepseekai2024deepseekv3technicalreport]. Here we assess the effectiveness of HMLA against hyperbolic Multi-Head self-Attention. We train HELM-MiCE with the same setup, where we replace the HMLA layers with a hyperbolic Multi-Head self-Attention layer as given in Equation˜3. We denote the this model as MiCE-HMHA. The results are shown in Table˜7. HELM-MiCE outperforms MiCE-HMHA in 3 out of the 5 tasks, achieving the same accuracy for 1 task, with the MiCE-HMHA achieving better accuracy in the last task. The results demonstrate the effectiveness of HELM-MiCE while significantly reducing the memory footprint of the KV-cache. Both hyperbolic models still outperform the Euclidean model, demonstrating the effectiveness of HELM in general.
D.4 Ablation for HoPE
In this section we assess the effectiveness of HoPE against other hyperbolic positional encoding methods, namely the learned relative positional encoding from Hypformer [yang2024hypformer]. We devise a variant of HELM-D, denoted as HELM-D-L and a variant of HELM-MiCE, denoted as HELM-MiCE-L, where each model uses the learned positional encoding instead of HoPE. The results are shown in Table˜8. Overall both HELM-MiCE and HELM-D outperform their counterparts that use learned positional encoding instead of HoPE. Interestly, however, HELM-MiCE-L and HELM-D-L outperformed HELM-MiCE and HELM-D respectively on the ARC-Challenging benchmark, possibly due to better alignment with reasoning prompts with non-uniform encodings. Nevertheless, the results demonstrate the effectiveness of HoPE over learned positional encodings in 4 out of the 5 tasks.
Table 9: Multi-choice answer accuracy for HELM-MiCE 1B and DeepSeekV3 1B across training stages. HELM-MiCE demonstrates consistent improvement over its Euclidean counter part.
Model Token Count CommonsenseQA HellaSwag OpenbookQA MMLU ARC-Challenging Avg
DeepseekV3 (1B) 4B 18.8 25.1 26.4 23.5 22.1 23.2
HELM-MiCE (1B) $\mathbf{4\mathrm{B}}$ $\mathbf{19.5}$ $\mathbf{26.0}$ $\mathbf{27.0}$ $\mathbf{25.6}$ $\mathbf{22.9}$ $\mathbf{24.3}$
DeepseekV3 (1B) 4.5B 19.0 26.2 27.2 23.6 22.6 23.7
HELM-MiCE (1B) $\mathbf{4.5\mathrm{B}}$ $\mathbf{19.7}$ $\mathbf{26.4}$ $\mathbf{27.6}$ $\mathbf{25.5}$ $\mathbf{23.1}$ $\mathbf{24.5}$
DeepseekV3 (1B) 5B 19.5 26.2 27.4 23.6 22.7 23.9
HELM-MiCE (1B) $\mathbf{5\mathrm{B}}$ $\mathbf{19.8}$ $\mathbf{26.5}$ $\mathbf{28.4}$ $\mathbf{25.9}$ $\mathbf{23.7}$ $\mathbf{24.9}$
D.5 Performance over Training Stages
While we don’t provide standard deviation from multiple runs for the 1B models as is typical for model of this size, we provide the model performance of HELM-MiCE and DeepSeekV3 1B in Table˜9. HELM-MiCE demonstrates consistent improvement over its Euclidean counter part, suggesting this improvement could sustain when the training corpus scales.